A chemist needs 18 liters of a 30% acid solution. How much of a 20% solution should be mixed with a 40% solution to get that (round your answer to the nearest liter)?
anyone who can show me how to do this great :) i was stuck on it.

Answer by stanbon(75887)

A chemist needs 18 liters of a 30% acid solution. How much of a 20% solution should be mixed with a 40%
-------
Equation:
acid + acid = acid
0.20x + 0.40(18-x) = 0.30*18
----
Multiply thru by 100 to get:
20x + 40*18 - 40x = 30*18
-20x = -10*18
x = 9 liters (amt. of 20% solution needed)
----
18-x = 9 liters (amt. of 40% solution needed)
=======================
Cheers,
Stan H.
====================